Explain the process and show EACH step in the process of simplifying the following problems:
b. sqrt(98c^3)
c. sqrt[200r^5(s^3)]

Answer by fractalier(6550)

In each of these, the idea is to factor the radicands into a perfect square and another factor...thus
sqrt(98c^3) = sqrt(49c^2) * sqrt(2c) = 7c * sqrt(2c)
and
sqrt[200r^5(s^3)] = sqrt(100(r^4)(s^2)) * sqrt(2rs) = 10(r^2)s * sqrt(2rs)